# Prospects for High-Frequency Gravitational-Wave Detection with GEO600

### v1

@@ I Introduction
Recent years have yielded a diverse set of proposals to improve the high frequency sensitivity of GW detectors. The LIGO and Virgo interferometers have developed both frequency-independent and frequency-dependent squeezing technologies, though in the O3, O4a and O4b observing runs, Virgo has only employed frequency-independent squeezing while LIGO has operated with frequency-dependent squeezing [9, 10, 11]. LIGO A+, Voyager, and A##\## are design concepts that aim to retrofit the existing LIGO observatory facilities by implementing upgraded technology [12, 13, 14]. Third generation ground interferometers [15, 16] will expand the sensitivity range of such experiments. At higher frequencies, resonant spheres [17, 18] could detect GWs up to $10$ kHz, optically-levitated sensors [19, 20] will search for GWs at tens to hundreds of kHz, and resonant electromagnetic detectors [21, 22, 23, 24, 25], bulk acoustic wave devices [26, 27, 28], and tabletop-scale interferometers [29, 30, 31] will probe frequencies at MHz and above. Additionally, space-based interferometers [32, 33, 34], atom interferometry experiments [35, 36, 37], and pulsar timing arrays (PTAs) [38, 39] will target frequencies below $1$ Hz. While such technologies for expanding the range of detectable GW frequencies are very promising — indeed, PTAs have recently yielded strong evidence for the existence of a stochastic GW background [40, 41] — the vast majority will require at least a decade of development to become operational.
@@
Detecting GWs at frequencies in the tens of kHz holds the potential to unveil signatures of new physics such as very light compact object binaries or ultralight boson clouds [42, 43]. Sub-solar mass (SSM) compact objects could be astrophysical objects like BHs created through an unconventional formation channel [44, 45, 46], a primordial BH [47, 48, 49], or some other exotic compact object [50, 51, 52, 53]. If multiple SSM objects form a binary, it will be detectable by searching for its effects with a high-frequency GW detector [54, 55]. Additionally, isolated spinning black holes can form boson clouds via the superradiance effect if there is some unknown ultralight boson cloud [56, 57]. Detection of the GWs emitted by the boson clouds could be evidence of new physics. The quantum chromodynamics (QCD) axion [58, 59, 60, 61] and the dark photon [62, 63, 64] are two prominent ultralight bosons that could have astrophysical effects in GWs [65, 66, 67]. The LVK collaboration has actively searched for GWs from SSM binaries [68, 69, 70, 71, 72] and superradiant boson clouds in their data [73, 74, 75]. SSM binaries and superradiant boson clouds may generate GWs with characteristic frequencies above the sensitivity bands of current generation detectors [76]. In addition to superradiance, ultralight bosons could be bound to the gravitational potential of a NS [77, 78] or BH which would allow indirect detection via fifth-force in GW signals [79, 80, 81] or in binary pulsar observations [82, 83, 84, 85]. Outside of GW observations, ultralight bosons and sub-solar mass compact objects have been searched for with both direct and indirect experiments; however, no evidence has been found [86, 87, 88, 89, 90, 91, 92, 93].
@@
The rest of this paper is organized as follows. In Sec. II, we discuss the underlying detector physics that motivated the investigation of GEO600 as a high-frequency GW detector. We describe our methods for computing strain sensitivities for detuned configurations of GEO600. We discuss the resulting sensitivities of different interferometer configurations and compare them to other ground-based interferometric detectors. In Sec. III, we introduce the high-frequency astrophysical sources and describe their GW waveforms and properties. We discuss how the SNR is calculated and the resulting ability of each interferometer to detect the sources. In Sec. IV, we conclude our work and discuss future research. Python code and Jupyter notebooks for generating the results shown in this document can be downloaded from GitHub [99].
@@ II.1 Underlying Physics
where $\nu$ is the laser frequency, and $n$ is an integer. For GEO600, $L_{\mathrm{SRC}}=L_{\mathrm{MI}}+L_{\mathrm{SR}}$ is the optical distance light travels from the MI’s end mirror to the MSR (when the detuning is zero). For this folded configuration, $L_{\mathrm{MI}}=2\times 600=1200$m. The definition of $L_{\mathrm{SRC}}$ ensures that $2\nu L_{\mathrm{SRC}}/c$ is an integer already; since GWs create sidebands at ±$\pm f_{\mathrm{GW}}$, the condition simplifies to one for the (positive) GW frequencies, which produce sidebands that are resonant in the SRC:
@@
[…]
@@ II.2 Methods
To construct the strain sensitivity function for a given configuration of GEO600, we include the following dominant noise sources $n_{i}$: laser quantum noise, classical laser amplitude and frequency noise, seismic noise, and thermal noise in the suspensions, substrates, and coatings of the optics. The expressions for the amplitude spectral densities for each noise source $A_{n_{i}}(f)$ (with units $[n_{i}]/\sqrt{\mathrm{Hz}}$) are presented in Appendix B.
@@
[…]
@@
Beam splitter (BS) 0.513872 $130$ ppm $9.3$ kg
Central East mirror (MCE) $13$ ppm $130$ ppm $5.6$ kg
Central North mirror (MCN) $13$ ppm $130$ ppm $5.6$ kg
Far East mirror (MFE) $8.3$ ppm $130$ ppm $5.6$ kg
Far North mirror (MFN) $8.3$ ppm $130$ ppm $5.6$ kg
Power-recycling mirror (MPR) $900$ ppm $130$ ppm $2.92$ kg
Signal-recycling mirror (MSR) $0.09995$ $50$ ppm $2.92$ kg
@@
where $\mathcal{R}(f)$ is the sky-averaged detector antenna pattern and $\left|F_{+}^{\mathrm{fold}}\left(f,\hat{n}=\hat{z}\right)\right|$ is the antenna pattern function evaluated for a GW incident from overhead. In the low-frequency limit $\mathcal{R}(f)=2/5$, but for frequencies comparable to the inverse light travel time in the Michelson arms, high frequency effects become important. In Appendix A, we derive the antenna pattern of a folded Michelson interferometer as done in Ref. [113] and clarify how the folded geometry modifies the high-frequency response.
@@
Note that when we plot detector strain sensitivities in this work (e.g., Figs. 2, 3, and 4), we assume a fiducial GW sky location of $\theta_{\hat{n}}=15$ and $\phi_{\hat{n}}=0$, as is typically done in the CE community [116, 117].
@@ III.1.1 GW Strain Model
The gravitational potential of a BH allows for particle bound states that can be approximated by hydrogenic wave functions with radial, orbital, and azimuthal quantum numbers $(n,l,m)$ [124]. Thus, the resulting boson cloud formed by superradiance can be described as a hydrogenic wave function, with all of the bosons occupying the same state. For our simulation, we only consider the fundamental radial quantum number $n=0$ and the dominant unstable mode $j=l+s=m=1$—where $l$, $s$, and $m$ are the orbital angular momentum, spin angular momentum, and magnetic quantum numbers, respectively. This unstable mode is chosen because it has the fastest superradiant growth rate in a given system [123, 125], making it the most dominant energy mode. Notably, this also requires that we assume that the BHs which source these scalar axion clouds are not primordial in origin, as previous works considering primordial BHs have found that only states with quantum numbers $n\geq 5$ can still exist and radiate GWs in the present day [42]. We furthermore constrain our BHs to have initial dimensionless spin parameter of $\chi_{\text{i}}=0.7$.
@@
$h_{0}^{v}\approx 1.28\times 10^{-22}\!\left(\!\frac{M}{0.3M_{\odot}}\!\right)\!\!\left(\frac{\alpha}{0.2}\right)^{\!5}\!\!\left(\!\frac{0.03\,\text{Mpc}}{r}\!\right)\!\!\left(\!\frac{0.7-\chi_{f}}{0.1}\!\right)\!,$ (7)
@@
$h_{0}^{s}\approx 10^{-25}\!\left(\!\frac{M}{0.3M_{\odot}}\!\right)\!\!\left(\frac{\alpha}{0.2}\right)^{\!7}\!\!\left(\!\frac{0.03\,\text{Mpc}}{r}\!\right)\!\!\left(\!\frac{0.7-\chi_{f}}{0.1}\!\right)\!,$ (10)
@@ III.2.1 GW Strain Model
Sub-solar mass compact objects are a compelling category of potential GW sources in the high-frequency regime accessible to current and future detectors. While conventional astrophysical compact objects such as neutron stars and black holes are typically formed with masses above $M_{\odot}$, various theoretical models predict the existence of sub-solar mass compact objects through unconventional formation channels or as entirely new types of exotic matter configurations [43]. Outside the conventional BH-formation theory of stellar evolution, PBHs may have formed from the gravitational collapse of density perturbations in the early Universe [49, 47, 48], and BHs formed from the gravitational collapse of dark matter halos [46] have been proposed as sub-solar mass compact objects. Additionally, there exist theories of sub-solar mass exotic compact objects, such as gravitino stars [52], boson stars [50], moduli stars [53], and gravastars [51]. If a sub-solar mass binary were found, classification between BBH and BNS is possible since tidal deformabilities strongly affect the waveform for light NS [126].
We simulate GWs from sub-solar mass compact binary mergers within the mass range of $10^{-6}~{}M_{\odot}$ to $1~{}M_{\odot}$. Note that the detection of GWs sourced from compact objects within this mass range would not be provide the sufficient parameters to distinguish that the object is a PBH specifically. In addition, if the source were a PBH, a detection within this mass range could only provide evidence that PBHs constitute some fraction of dark matter [76]. We compute the characteristic strain from the binary merger of sub-solar mass compact objects using the inspiral, merger, and ringdown waveforms available in PyCBC’s IMRPhenomD package [127]. For simplicity, we assume an inclination angle of $0\degree$, making $\tilde{h}_{+}(f)\!=\!\tilde{h}_{\times}(f)$ and $\tilde{h}(f)=\mathcal{R}(f)\tilde{h}_{0}(f)$ (see Appendix A.3). We calculated the sky-averaged SNR using the chirp formula given by [6] as
[…]
@@ III.2.2 Detection Prospects
Our figure of merit to compare the detection capabilities of our detuned GEO600 configurations against LIGO is the sky-averaged SNR produced by a characteristic high-frequency SSM compact binary merger in each detector. In Fig. 6, we show the GW strain of compact mergers with individual BH masses of $10^{-1}$, $10^{-2}$, and $10^{-3}~{}M_{\odot}$, located a distance of 30 kPc from the detector. As the chirp mass decreases, $f_{\text{min}}$ increases while the GW strain is reduced. The frequency-domain signal is calculated with an observation time of one week, and the SNR is related to the area in the plot between the strain and noise spectral functions. We find that there are not sub-solar mass binaries which can be detected easier with a detuned GEO600 detector. Over the frequency range where the detuning-scanned GEO600 has a better sensitivity than aLIGO, the strain from the merger is weak, and thus the accumulated SNR is so little that it does not sway the total SNR in favor of GEO600. While the time domain signal is stronger at high frequencies, the early inspiral dominates the SNR because it has more cycles in band. If there were a sub-solar mass binary that were very slowly evolving in frequency at a timescale of $\sim$ months, then detuning GEO to the exact frequency could improve the sensitivity compared with LIGO. For nearly monochromatic sub-solar mass signals, one would see a similar improvement as to what we saw in Sec. III.1. Thus, GEO600 is not an optimal ground-based interferometer for the detection of most sub-solar mass compact binary mergers.
@@ IV Conclusion
In this work, we investigated how the high-frequency sensitivity of GEO600 can be scanned by detuning the location of the MSR. The MSR creates an optical cavity that enables narrow-band resonant amplification of GW-driven sideband fields. Detuning the MSR changes which frequencies satisfy the resonant condition, allowing GEO600 to scan over a range of high frequencies with broadband sensitivity. We estimated the SNR in the experiment for characteristic high-frequency GW sources located within the Milky Way. This setup of detuning GEO600 does not result in an improvement in searches for sub-solar mass binaries due to their long inspiral in low frequencies, and the bandwidth being $\sim 2$ kHz. However, GEO600 could outperform the aLIGO design sensitivity for the detection of GWs sourced from nearly monochromatic sources like boson clouds within the galaxy. The interferometer would be able to detect GWs sourced from vector boson clouds at frequencies between $500$ Hz to $31.3$ kHz, gaining the advantage over the aLIGO design sensitivity at $15.1$ kHz and over O4 LIGO at $10$ kHz.

### v2

@@ I Introduction
Recent years have yielded a diverse set of proposals to improve the high frequency sensitivity of GW detectors. The LIGO and Virgo interferometers have developed both frequency-independent and frequency-dependent squeezing technologies, though in the O3, O4a and O4b observing runs, Virgo has only employed frequency-independent squeezing while LIGO has operated with frequency-dependent squeezing [9, 10, 11]. LIGO A+, Voyager, and A#\# are design concepts that aim to retrofit the existing LIGO observatory facilities by implementing upgraded technology [12, 13, 14]. Third generation ground interferometers [15, 16] will expand the sensitivity range of such experiments. At higher frequencies, resonant spheres [17, 18] could detect GWs up to $10$ kHz, optically-levitated sensors [19, 20] will search for GWs at tens to hundreds of kHz, and resonant electromagnetic detectors [21, 22, 23, 24, 25], bulk acoustic wave devices [26, 27, 28], and tabletop-scale interferometers [29, 30, 31] will probe frequencies at MHz and above. Additionally, space-based interferometers [32, 33, 34], atom interferometry experiments [35, 36, 37], and pulsar timing arrays (PTAs) [38, 39] will target frequencies below $1$ Hz. While such technologies for expanding the range of detectable GW frequencies are very promising — indeed, PTAs have recently yielded strong evidence for the existence of a stochastic GW background [40, 41] — the vast majority will require at least a decade of development to become operational.
@@
Detecting GWs at frequencies in the tens of kHz holds the potential to unveil signatures of new physics such as very light compact object binaries or ultralight boson clouds [42, 43]. Sub-solar mass (SSM) compact objects could be astrophysical objects like BHs created through an unconventional formation channel [44, 45, 46], a primordial BH [47, 48, 49], or some other exotic compact object [50, 51, 52, 53]. If multiple SSM objects form a binary, it will be detectable by searching for its effects with a high-frequency GW detector [54, 55]. Additionally, isolated spinning black holes can form boson clouds via the superradiance effect if there is some unknown ultralight boson cloud [56, 57]. Detection of the GWs emitted by the boson clouds could be evidence of new physics. The quantum chromodynamics (QCD) axion [58, 59, 60, 61] and the dark photon [62, 63, 64] are two prominent ultralight bosons that could have astrophysical effects in GWs [65, 66, 67]. The LVK collaboration has actively searched for GWs from SSM binaries [68, 69, 70, 71, 72] and superradiant boson clouds in their data [73, 74, 75]. SSM binaries and superradiant boson clouds may generate GWs with characteristic frequencies above the sensitivity bands of current generation detectors [76]. In addition to superradiance, ultralight bosons could be bound to the gravitational potential of a NS [77, 78] or BH which would allow indirect detection via fifth-force in GW signals [79, 80, 81] or in binary pulsar observations [82, 83, 84, 85]. Outside of GW observations, ultralight bosons and SSM compact objects have been searched for with both direct and indirect experiments; however, no evidence has been found [86, 87, 88, 89, 90, 91, 92, 93]. Lastly, enhancing the high-frequency sensitivity of gravitational-wave detectors expands the range of astrophysical phenomena that can be probed through data analysis [94].
@@
The rest of this paper is organized as follows. In Sec. II, we discuss the underlying detector physics that motivated the investigation of GEO600 as a high-frequency GW detector. We describe our methods for computing strain sensitivities for detuned configurations of GEO600. We discuss the resulting sensitivities of different interferometer configurations and compare them to other ground-based interferometric detectors. In Sec. III, we introduce the high-frequency astrophysical sources and describe their GW waveforms and properties. We discuss how the SNR is calculated and the resulting ability of each interferometer to detect the sources. In Sec. IV, we conclude our work and discuss future research. The data used in this paper can be downloaded from GitHub [100].
@@ II.1 Underlying Physics
where $\nu$ is the laser frequency, and $n$ is an integer. For GEO600, $L_{\mathrm{SRC}}=L_{\mathrm{MI}}+L_{\mathrm{SR}}$ is the optical distance light travels from the MI’s end mirror to the MSR (when the detuning is zero). For this folded configuration, $L_{\mathrm{MI}}=2\times 600=1200$m. The definition of $L_{\mathrm{SRC}}$ ensures that $2\nu L_{\mathrm{SRC}}/c$ is an integer already; since GWs create sidebands at $\pm f_{\mathrm{GW}}$, the condition simplifies to one for the (positive) GW frequencies, which produce sidebands that are resonant in the SRC:
@@
For this work, we primarily target the frequency which appears for n=0n=0111The resonant frequencies for $|n|>0$ are less advantageous for detection because the GW-induced phase delay in the MI is diminished for frequencies greater than $f_{\mathrm{fsr}}=c/2L_{\mathrm{src}}$ (see Appendix A).. The detuning angle $\phi$ is related to the MSR displacement $\delta l$ by $\phi=k\delta l$, where $k=2\pi\nu/c$. Therefore, by finely tuning the MSR location, we can scan the SRC resonant frequencies, and thus the resonant peaks in the GEO600 detector response, over a broad frequency range [106]. Detuning the MSR allows for lowest-order resonant peaks in the detector response function to occur as high as 62.5 kHz.
@@ II.2 Methods
To construct the strain sensitivity function for a given configuration of GEO600, we include the following dominant noise sources $n_{i}$: laser quantum noise, classical laser amplitude and frequency noise, seismic noise, and thermal noise in the suspensions, substrates, and coatings of the optics. The expressions for the amplitude spectral densities for each noise source $A_{n_{i}}(f)$ (with units $n_{i}]/\sqrt{\mathrm{Hz}}$) are presented in Appendix B.
@@
where the summation occurs over all the noise sources $n_{i}$, $S_{n_{i}}\equiv A_{n_{i}}^{2}$ is the PSD of the noise source $n_{i}$, and $\partial P/\partial n_{i}$ is the transfer function from noise source $n_{i}$ to the detection observable $P$. In this paper, we designate the observable as the power at the dark port of the interferometer — see the optical layout depicted in Fig. 1. Lastly, $\langle\partial P/\partial h\rangle$ is the transfer function from an incident GW to the observable (i.e., the detector response function). We use the notation $\langle\cdot\rangle$ to represent a sky-averaged quantity. To summarize, the transfer function $\partial P/\partial n_{i}$ projects a given noise source onto the readout, and $\partial P/\partial h$ projects the GW strain onto the same readout; thus, Eq. (3) projects each noise source into some equivalent sky-averaged strain noise.
@@
Beam splitter
$0.513872$ $130$ ppm $9.3$ kg
Central East mirror
$13$ ppm $130$ ppm $5.6$ kg
Central North mirror
$13$ ppm $130$ ppm $5.6$ kg
Far East mirror
$8.3$ ppm $130$ ppm $5.6$ kg
Far North mirror
$8.3$ ppm $130$ ppm $5.6$ kg
Power-recycling mirror
$900$ ppm $130$ ppm $2.92$ kg
Signal-recycling mirror
$0.09995$ $50$ ppm $2.92$ kg
@@
[…]
@@
Note that when we plot detector strain sensitivities in this work (e.g., Figs. 2, 3, 4, and 6), we assume a fiducial GW sky location of $\theta_{\hat{n}}=15$ and $\phi_{\hat{n}}=0$, as is typically done in the CE community [117, 118].
@@ III.1.1 GW Strain Model
The gravitational potential of a BH allows for particle bound states that can be approximated by hydrogenic wave functions with radial, orbital, and azimuthal quantum numbers $(n,l,m)$ [125]. Thus, the resulting boson cloud formed by superradiance can be described as a hydrogenic wave function, with all of the bosons occupying the same state. For our simulation, we only consider the fundamental radial quantum number $n=0$ and the dominant unstable mode j=l+s=m=1j=l+s=m=1—where $l$, $s$, and $m$ are the orbital angular momentum, spin angular momentum, and magnetic quantum numbers, respectively. This unstable mode is chosen because it has the fastest superradiant growth rate in a given system [124, 126], making it the most dominant energy mode. Notably, this also requires that we assume that the BHs which source these scalar axion clouds are not primordial in origin, as previous works considering primordial BHs have found that only states with quantum numbers $n\geq 5$ can still exist and radiate GWs in the present day [42]. We furthermore constrain our BHs to have initial dimensionless spin parameter of $\chi_{\text{i}}=0.7$.
@@
$\hskip-7.59995pth_{0}^{v}\approx 1.28\times 10^{-22}\!\left(\!\frac{M}{0.3M_{\odot}}\!\right)\!\!\left(\frac{\alpha}{0.2}\right)^{\!5}\!\!\left(\!\frac{0.03\,\text{Mpc}}{r}\!\right)\!\!\left(\!\frac{0.7-\chi_{f}}{0.1}\!\right)\!,$ (7)
@@
$\hskip-8.99994pth_{0}^{s}\approx 10^{-25}\!\left(\!\frac{M}{0.3M_{\odot}}\!\right)\!\!\left(\frac{\alpha}{0.2}\right)^{\!7}\!\!\left(\!\frac{0.03\,\text{Mpc}}{r}\!\right)\!\!\left(\!\frac{0.7-\chi_{f}}{0.1}\!\right)\!,$ (10)
@@ III.2.1 GW Strain Model
Sub-solar mass (SSM) compact objects are a compelling category of potential GW sources in the high-frequency regime accessible to current and future detectors. While conventional astrophysical compact objects such as neutron stars and black holes are typically formed with masses above $M_{\odot}$, various theoretical models predict the existence of SSM compact objects through unconventional formation channels or as entirely new types of exotic matter configurations [43]. Outside the conventional BH-formation theory of stellar evolution, PBHs may have formed from the gravitational collapse of density perturbations in the early Universe [49, 47, 48], and BHs formed from the gravitational collapse of dark matter halos [46] have been proposed as SSM compact objects. Additionally, there exist theories of SSM exotic compact objects, such as gravitino stars [52], boson stars [50], moduli stars [53], and gravastars [51]. If a SSM binary were found, classification between BBH and BNS is possible since tidal deformabilities strongly affect the waveform for light NS [127].
We simulate GWs from SSM compact binary mergers within the mass range of $10^{-6}~M_{\odot}$ to $1~M_{\odot}$. Note that the detection of GWs sourced from compact objects within this mass range would not be provide the sufficient parameters to distinguish that the object is a PBH specifically. In addition, if the source were a PBH, a detection within this mass range could only provide evidence that PBHs constitute some fraction of dark matter [76]. We compute the characteristic strain from the binary merger of SSM compact objects using the inspiral, merger, and ringdown waveforms available in PyCBC’s IMRPhenomD package [128]. For simplicity, we assume an inclination angle of $0\degree$, making $\tilde{h}_{+}(f)\!=\!\tilde{h}_{\times}(f)$ and $\tilde{h}(f)=\mathcal{R}(f)\tilde{h}_{0}(f)$ (see Appendix A.3). We calculated the sky-averaged SNR using the chirp formula given by [6] as
$\langle\rho^{2}\rangle=4\int_{f_{\text{min}}}^{\infty}\frac{|\tilde{h}(f)|^{2}}{\langle S_{h}(f)\rangle}df,$ (13)
@@ III.2.2 Detection Prospects
Our figure of merit to compare the detection capabilities of our detuned GEO600 configurations against LIGO is the sky-averaged SNR produced by a characteristic high-frequency SSM compact binary merger in each detector. In Fig. 6, we show the GW strain of compact mergers with individual BH masses of $10^{-1}$, $10^{-2}$, and $10^{-3}~M_{\odot}$, located a distance of 30 kPc from the detector. As the chirp mass decreases, $f_{\text{min}}$ increases while the GW strain is reduced. The frequency-domain signal is calculated with an observation time of one week, and the SNR is related to the area in the plot between the strain and noise spectral functions. We find that there are not SSM binaries which can be detected easier with a detuned GEO600 detector. Over the frequency range where the detuning-scanned GEO600 has a better sensitivity than aLIGO, the strain from the merger is weak, and thus the accumulated SNR is so little that it does not sway the total SNR in favor of GEO600. While the time domain signal is stronger at high frequencies, the early inspiral dominates the SNR because it has more cycles in band. If there were a SSM binary that were very slowly evolving in frequency at a timescale of $\sim$ months, then detuning GEO to the exact frequency could improve the sensitivity compared with LIGO. For nearly monochromatic SSM signals, one would see a similar improvement as to what we saw in Sec. III.1. Thus, GEO600 is not an optimal ground-based interferometer for the detection of most SSM compact binary mergers.
@@ IV Conclusion
In this work, we investigated how the high-frequency sensitivity of GEO600 can be scanned by detuning the location of the MSR. The MSR creates an optical cavity that enables narrow-band resonant amplification of GW-driven sideband fields. Detuning the MSR changes which frequencies satisfy the resonant condition, allowing GEO600 to scan over a range of high frequencies with broadband sensitivity. We estimated the SNR in the experiment for characteristic high-frequency GW sources located within the Milky Way. This setup of detuning GEO600 does not result in an improvement in searches for SSM binaries due to their long inspiral in low frequencies, and the bandwidth being $\sim 2$ kHz. However, GEO600 could outperform the aLIGO design sensitivity for the detection of GWs sourced from nearly monochromatic sources like boson clouds within the galaxy. The interferometer would be able to detect GWs sourced from vector boson clouds at frequencies between $500$ Hz to $31.3$ kHz, gaining the advantage over the aLIGO design sensitivity at $15.1$ kHz and over O4 LIGO at $10$ kHz.